On 1/1/22 Big Co acquired 100% of Little Co voting stock for $500,000. On that date Little's book value was $350,000. Little had the following misvalued assets and liabilities at 1/1/22:

Land, undervalued by $20,000

Inventory, FIFO basis, undervalued by $10,000

PPE, 5 year remaining life, undervalued by $80,000.

Bonds payable, 10 year life, undervalued by $20,000

During 2022 Little reported earnings of $80,000 and paid dividends of $10,000

1. How much goodwill was inherent in the purchase of Little?

2. Assuming Big uses the equity method, what would be the reported investment income for 2022?

3. What would be the balance in the "Investment in Little" after all of the 2022 entries?
Answer #1

ANSWER

2) Amount $
Little earnings        80,000
Less: Inventory undervalued FIFO        10,000
Less: PPE Amortization        16,000 =80000/5
Add: Bond Payable Undervalued          2,000 =20000/10
Adjusted earnings of of Little        56,000
Reported Investment income for 2022 $ 56,000 =56000*100%
3) Amount $
Purchase Cost     500,000
Add: Investment income for 2022 56,000
Less: Dividend received ( 10,000 x 100% ) -10,000
Balance in Investment in little    546,000

================

DEAR STUDENT,

IF YOU HAVE ANY QUERY PLEASE ASK ME IN THE COMMENT BOX,I AM HERE TO HELP YOU.PLEASE GIVE ME POSITIVE RATING..

****************THANK YOU****************